OpenStudy (ladylauren):

Write one example of a linear function and one example of a non-linear function. (Use x and y as the variables)

OpenStudy (dayakar):

f(x,y)=ax+by+c [ linear function] example f(x,y) = 3x +y +8

OpenStudy (dayakar):

non- linear function f(x,y) = x^2 - 3xy +y^2

OpenStudy (ladylauren):

that doesn't make sense i thought liner functions were like (1,2), (2,3), (3,4)

OpenStudy (ladylauren):

in a straight line

OpenStudy (ladylauren):

i just don't know how to write it with x and y

OpenStudy (dayakar):

linear equation is in the form of ax +by +c=0 3x+5y+9=0

OpenStudy (dayakar):

it shows a straight line

OpenStudy (ladylauren):

ok

OpenStudy (dayakar):

is it clear

OpenStudy (ladylauren):

kinda